OpenStudy (anonymous):

What is the common ratio of the geometric sequence whose first term is 27 and fourth term is 64? (1) 3/4 (3) 4/3 (3) 64/81 (4) 37/3

OpenStudy (anonymous):

Ok so in a geometric sequence, remember that each term is something multiplied by the last term, and that this something doesn't change. If the first term is 27, then \[27*x\] gives the second term, \[27*x*x\] gives the third term, \[27*x*x*x\] gives the fourth term. Thus, \[27*x*x*x=64\] \[x^3=64/27\] \[x=4/3\]

OpenStudy (anonymous):

okay great thanks!